OpenStudy (anonymous):

What is the equation of the parabola with a focus at (0, -5) and directrix y = 5?

OpenStudy (anonymous):

To derive a quadratic equation using distances, used the combined distance formula. The combined distance formula is \[\sqrt{(y_2-y_1)^2+(y_2-y_1)^2}=\sqrt{(y_2-y_1)^2}\]. Plug in the focus on the left side of the equation, and plug the directrix on the right side.

OpenStudy (anonymous):

y=4py directix=-p

OpenStudy (anonymous):

y^2=4px

OpenStudy (anonymous):

of course you have to find out which side it opens to